OpenStudy (anonymous):

what is the slope ,x-intercept and y intercept of 3x+y=7

OpenStudy (anonymous):

y = mx + b , where m is the slope and b is the y intercept

OpenStudy (anonymous):

The y intercept is where x=0

OpenStudy (anonymous):

x intercept is where y=0

OpenStudy (anonymous):

slope -3

OpenStudy (anonymous):

x intercept = 7/3 & y intercept = 7

OpenStudy (anonymous):

correct